Using the data below, perform the following tasks


23 26 29 32 35



A. Construct a sampling distribution of the mean when n=3



B. Draw a histogram for the sample means

1
Expert's answer
2022-06-06T23:11:54-0400

A.

We have population values 23, 26, 29, 32, 35, population size N=5 and sample size n=3.

Mean of population "(\mu)" = "\dfrac{23+26+29+32+35}{5}=29"

Variance of population 


"\sigma^2=\dfrac{\Sigma(x_i-\bar{x})^2}{N}=\dfrac{36+9+0+9+36}{5}=18"




"\sigma=\sqrt{\sigma^2}=\sqrt{18}=3\sqrt{2}\approx4.24264"


The number of possible samples which can be drawn without replacement is "^{N}C_n=^{5}C_3=10."

"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 no & Sample & Sample \\
& & mean\ (\bar{x})
\\ \hline
 1 & 23,26,29 & 26 \\
 \hdashline
 2 & 23,26,32 & 27 \\
 \hdashline
 3 & 23,26,35 & 28\\
 \hdashline
 4 & 23,29,32 & 28 \\
 \hdashline
 5 & 23,29,35 & 29 \\
 \hdashline
 6 & 23,32,35 & 30 \\
 \hdashline
 7 & 26,29,32 & 29 \\
 \hdashline
 8 & 26,29,35 & 30 \\
 \hdashline
 9 & 26,32,35 & 31 \\
 \hdashline
 10 & 29,32,35 & 32 \\
 \hdashline
\end{array}"




"\def\arraystretch{1.5}
 \begin{array}{c:c:c:c:c}
 \bar{X} & f(\bar{X}) &\bar{X} f(\bar{X})& \bar{X}^2f(\bar{X}) 
\\ \hline
26 & 1/10 & 26/10 & 676/10 \\
 \hdashline
 27 & 1/10 & 27/10 & 729/10 \\
 \hdashline
 28 & 2/10 & 56/10 & 1568/10 \\
 \hdashline
 29 & 2/10 & 58/10 & 1682/10 \\
 \hdashline
 30 & 2/10 & 60/10 & 1800/10 \\
 \hdashline
 31 & 1/10 & 31/10 & 961/10 \\
 \hdashline
 32 & 1/10 & 32/10 & 1024/10 \\
 \hdashline
\end{array}"



Mean of sampling distribution 


"\mu_{\bar{X}}=E(\bar{X})=\sum\bar{X}_if(\bar{X}_i)=29=\mu"



The variance of sampling distribution 


"Var(\bar{X})=\sigma^2_{\bar{X}}=\sum\bar{X}_i^2f(\bar{X}_i)-\big[\sum\bar{X}_if(\bar{X}_i)\big]^2"




"=\dfrac{8440}{10}-(29)^2=3= \dfrac{\sigma^2}{n}(\dfrac{N-n}{N-1})"




"\sigma_{\bar{X}}=\sqrt{3}\approx1.73205"



B.